Sol Summary – May 26th
Crew 281 Sol Summary Report 26 May 2023
Sol: 12
Summary Title: The bittersweet ending
Author’s name: Ritupriya Patil (XO)
Mission Status: Nominal
Sol Activity Summary:
EVA:
Today was team’s final and longest, all female EVA. 4 hours of walking, climbing (a little) and driving the rover with the space suits and life support systems on their backs. The high dessert “Martian” environment reminded the crew that space is challenging. However, the crew forged through with their love for exploration. The three-part objective of today’s EVA was to run payload drone tests, acquire images for previously done geotechnical survey, and explore the Special Region (aka Hanksville Burpee Quarry).
The crew at 0738 hrs, finished the rover checks, then continued to set up for a drone payload test. The objective was to deliver a lightweight medical kit to a simulated injured astronaut. The experimental drone flights were to establish a proof of concept in an emergency scenario, and in extreme climate (like a high desert). MDRS was the perfect place for this. 9 tests were performed, 3 out of those were with zero payload to set up a baseline experiment for a DJI Mavic mini weighing about 250 g. In honor of Team Pegasus, the drone was affectionately named “Peggy”. Ana (Crew Scientist) enacted as the injured analogue astronaut, KC (Crew Engineer) was the HabCom and recorded the flight data; Ritu (XO, EVA Lead) was the drone pilot and experimental flight test lead. There were 2 payloads 25 g and 40 g that were retrieved by Ana, stationed 60 m away from the pilot during the 6 payload tests. The crew was successful in procuring the payload and flying the drone back safely with the harness intact. Special care was taken, and wind checks were performed prior to each flight. Post flight battery usage after each flight varied based on the combination of the payload weight and winds (between 3-5 knots). On an average the battery usage was from 8% – 15% on the payload flights (higher usage for heavier payload).
The data from this experiment will be used in calculating and scaling the analysis for a bigger drone system that would be capable of carrying heavier load and be flight capable in high winds, and lower air density to simulate the Martian atmosphere. This analysis will further feed into design to build a drone delivery system on Mars.
Next stop for the crew was Marble Ritual, to acquire vertical images and videos for previously done geotechnical survey at the site. Some great shots of the MDRS campus tucked in the red “Martian” terrain were captured from Marble Ritual.
Final objective for today’s EVA was to explore the Special Region. The crew drove north on Cow Dung Road passing the now known landmarks of Pooh’s Corner, Galileo Road, Cowboy Corner, intersection to Brahe Hwy 1572, and then left on Gateway to Lith to reach the destination, formerly known as Burpee Quarry. The crew witnessed sediment depositions from a large river that must have run through the quarry area and created a bend. Ritu got aerial imagery of the location and Ana collected 7 rock samples (226g) and soil samples (299g). The crew accomplished all the EVA goals for the last one for this mission; and with a bittersweet emotion drove back to the hab.
For lunch the crew had leftovers, also Megan (Commander) baked brownies to celebrate the crew’s final EVA accomplishment. After munching on the delicious treat, Megan, Ritu, and Ana helped with the GreenHab harvest and clean up. KC completed inventory lists for the end of mission checks. It was a long day and no time to rest since the simulation would end at 6 pm. We also squeezed in time for report and mission summary writing.
At 6 pm the crew opened the Hab airlock and stepped out in their flight suits, they were also wearing the mission patches and respective country flags (India, Portugal and USA) with great pride! Sergii (Mission Support) captured the end of mission walk out of the airlock. This was end of the 12 Sols we spent in the Hab: working on 13 EVAs, running the Ham Radio for 6 days, keeping the GreenHab operational, conducting 4 geotechnical surveys of the area, recording drone footage on all EVAs and conducting 19 drone payload tests, and other experiments. Sergii was invited for dinner after the simulation ended and everyone enjoyed the salad and herbs harvested from the GreenHab; along with Rajma Chawal (Indian red bean stew served with rice) that Ritu made for dinner. A nice cup of cardamom spiced chai was a perfect end for the crew’s last dinner together in the hab.
The crew did a sweep through of all the buildings and started wrapping up the experiments and cleaning. Some walked to the Musk Observatory for star gazing and enjoyed the last starry night at MDRS.
It was an honor to be able to participate in exploration of this “Martian” landscape. This was a humbling experience, and the crew would do it again in a heartbeat!
Look Ahead Plan: End of mission. The crew will exit the hab at noon tomorrow.
Anomalies in work: None
Weather: Winds
Crew Physical Status: Nominal
EVA: Completed 1 today
Reports to be filed:
● Greenhab Report
● EVA Report
● Ops Report
● Journalist Report
● PHOTOS!
● Photo of the Day
Support Requested: None
GreenHab Report – May 26th
Crew 281 GreenHab Report 26 May 2023
GreenHab Officer: Megan Kane
Environmental control: NA
Average temperatures: NA
Hours of supplemental light: None
Daily water usage for crops 0 gallons
Daily water usage for research and/or other purposes: 1 gallon
Water in Blue Tank 152.59 gallons
Time(s) of watering for crops:
NA
Changes to crops: None
Narrative:
As this is the final day of the season the greenhab was opened in the morning. Everything edible was harvested. Then the biomass was placed into trash bags. Once the primary biomass was removed the roots were removed from the planters. Some small roots remained in the large planters and soil. The individual planters had their soil deposited into the red wheelbarrow. The large blue buckets kept the soil that was in them. We documented this process.
After finishing with the plants, the equipment was wiped down and stored. The floor swept and the tables wiped down.
Harvest: (include which crop and mass in grams):
Tomatoes 3668
Lettuce 651
Micro greens 59
Chives 76
sage 42
basil 55
cucumber 1647
Support/supplies needed: None
EVA Report – May 26t
Crew 281 EVA Report 26-05-2023
EVA # 13
Authors: Ritupriya Patil (Executive Officer, EVA Lead), Ana Pires (Crew Scientist)
Purpose of EVA:
Drone flight to record and do aerial imagery for new exploration areas
Acquire vertical images and videos for previously done geotechnical survey of Marble Ritual (Not completed)
Start time: 0738 hrs
End time: 1123 hrs
Narrative:
Today was the longest EVA a crew of two. At 0738 hrs the crew left the airlock for rover checks, then continued to set up for a drone payload test. The objective was to deliver a lightweight medical kit to a simulated injured astronaut, this was to establish a proof of concept for flying drone in an emergency scenario in extreme climate (essentially a high desert); and MDRS was the perfect place. 9 tests were performed, 3 out of those were with zero payload to set up a baseline for the experiment. Ana (Crew Scientist) enacted as the injured analogue astronaut, KC (Crew Engineer) was the HabCom and recorded the flight data; Ritu (XO, EVA Lead) was the drone pilot and experimental flight test lead. There were 2 payloads 25 g and 40 g that were retrieved by Ana, stationed 60 m away from the pilot during the 6 tests. The crew was successful in procuring the payload and flying the drone back safely with the harness intact. Special care was taken, and wind checks were performed prior to each flight. Post flight battery usage after each flight varied based on the combination of the payload weight and winds (between 3-5 knots). On an average the battery usage was from 8% – 15% on the payload flights (higher usage for heavier payload).
The data from this experiment will be used in calculating and scaling the analysis for a bigger drone system that would be capable of carrying heavier load and be flight capable in high winds, and lower air density. This analysis will further feed into design concepts to aid to drone (or a similar system) delivery on Mars.
Next stop for the crew was Marble Ritual, to acquire vertical images and videos for previously done geotechnical survey at the site. Some great shots of the MDRS campus tucked in the red “Martian” terrain were captured from Marble Ritual.
Final objective for today’s EVA was to explore the Special Region. The crew drove north on Cow Dung Road passing the now known landmarks of Pooh’s Corner, Galileo Road, Cowboy Corner, intersection to Brahe Hwy 1572, and then left on Gateway to Lith to reach the destination, formerly known as Burpee Quarry. The crew witnessed sediment depositions from a large river that must have run through the quarry area and created a bend. Ritu got aerial imagery of the location and Ana collected 7 rock samples (226g) and soil samples (299g). The crew accomplished all the EVA goals and with a bittersweet emotion drove back to the hab after their final EVA for the mission. It was an honour to be able to participate in exploration of this “Martian” landscape. It was a humbling experience and the crew would do it again in a heartbeat!
Overview of EVA
Drone payload test to simulate medical emergency delivery
Drive on Cow Dung Road to the North, until Marble Ritual, park and walk right hand side to Marble Ritual scanline boulders
Drive on Cow Dung road further North, take left on Gateway to Lith. Drive 200 m, park on the side of the road and walk no more than 500 m radius
Issues Experienced During EVA
Intermittent comms issues.
Outcomes
Crew completed Rover checks
Acquired vertical images and videos for previously done geotechnical survey of Marble Ritual
Explored Special region, collected rock and soil samples and drone footage
Collected 7 rocks, (226g) and soil samples (299g)
Destination:
MDRS Hab and surrounding 500 m radius
Marble Ritual, walking no more than a radius of 500m;
Special Region
Coordinates :
(use UTM WSG 84): 518100 E, 4250900 N
(use UTM WSG 84): 518687 E, 4250789 N
(use UTM WSG 84): 518424E, 4256117N
EVA Participants 2: Ritupriya Patil (XO, EVA Lead), Ana Pires (Crew Scientist)
Road(s) and routes per MDRS Map:
Walk around the Hab
Drive on Cow Dung Road to the North, until Marble Ritual, park and walk right hand side to Marble Ritual scanline boulders
Drive on Cow Dung road further North, take left on Gateway to Lith. Drive 200 m, park on the side of the road and walk no more than 500 m radius
Mode of travel: Rover
Operations Report – May 26th
Crew 281 Operations Report 26 May 2023
SOL: 12
Name of person filing report: Megan Kane (Commander), Ritupriya Patil (Executive Officer) and KC Shasteen (Crew Engineer)
Non-nominal systems:
Electrical outlets in kitchen
Remote observatory
EVA Helmet
Notes on non-nominal systems: None
ROVERS
Spirit rover used: No
Hours: (before EVA): 222.5
Hours: (after EVA): NA
Beginning charge: (Before EVA): 100%
Ending charge: (On return from EVA, before recharging): NA
Currently charging: Yes
Opportunity rover used: Yes
Hours: (before EVA): 132.1
Hours: (after EVA): 133.2
Beginning charge: (Before EVA): 100%
Ending charge: (On return from EVA, before recharging): 82%
Currently charging: Yes
Curiosity rover used: No
Hours: (before EVA): 241.0
Hours: (after EVA): NA
Beginning charge: (Before EVA): 100%
Ending charge: (On return from EVA, before recharging): NA
Currently charging: Yes
Perseverance rover used: No
Hours: (before EVA): 264.6
Hours: (after EVA): NA
Beginning charge: (Before EVA): 100%
Ending charge: (On return from EVA, before recharging): NA
Currently charging: Yes
General notes on rovers: Oppy used on long EVA to the special region north of the Gateway to Lith. Operation was nominal.
Summary of Hab operations:
WATER USE: 22.1 gallons
Water (static tank): 180.0 gallons
Static tank pipe heater (on or off): off
Static tank heater (On or off): off
Toilet tank emptied: no
Summary of internet:
nominal
Summary of suits and radios: All suits and radios are fully recharged.
Summary of GreenHab operations:
WATER USE: 1 gallons
Heater: Off
Supplemental light: Off
Harvest: in grams :
Tomatoes 3668
Lettuce 651
Micro greens 59
Chives 76
sage 42
basil 55
cucumber 1647
Summary of Science Dome operations:
Dual split: Heat and AC, Off
Cacao Experiment:
Watered cacao seedlings
Summary of RAM operations: (Tools used, work done)
None
Summary of any observatory issues: NA
Summary of health and safety issues:
No new issues to report.
Questions, concerns and requests to Mission Support: None
Mission Summary – May 26th
Crew 281 Mission Summary
Date: 05/26/2023
As the final MDRS Crew of the season, Team Pegasus (Crew 281) finished the season strong:
doing 13 EVAs, running the Ham Radio for 6 days including the first Morse code contact, kept
the GreenHab operational, conducted 4 geotechnical surveys of the area, 19 drone payload tests
and other experiments. The five person team successfully maintained the Hab and conducted
operations for all 12 sols of the mission. Here is the team:
● Megan Kane: Commander and GreenHab Officer
● Ritupriya (Ritu) Patil: Executive Officer
● Rachel Jones: Health Safety Officer
● Ana Pires: Crew Scientist
● KC Shasteen: Crew Engineer
The 12 Sol mission was packed with research, testing and operational tasks. The team did take
the time to enjoy the mission and time together with good food from the three cooks: KC, Megan
and Ritu.
Mars Desert Research Station Facilities:
The team utilized most of the facilities over the course of the mission.
The Science Dome was home to several experiments. In the grow tent, 75 Cacao (Chocolate) seeds of 2 varieties were sprouted. Started on Sol 1, by Sol 12 the cotyledons had begun to emerge. The workspace was used to test and prepare the equipment for the geotechnical surveys and examine the samples. Additionally, the team supported maintaining the algae experiment from Crew 261. Lastly, Team Pegasus recorded several STEM educational videos there.
In the GreenHab, the commander and crew engineer worked to keep the plants alive late into the increasing heat. Both having experience with greenhouses and vertical farms, they worked to decrease the rising temperatures and increase the humidity for the plants. This was met with limited success due to availability of materials. The watering test conducted by the commander went well.
The Repair and Maintenance (RAM) module was used to support many activities including drone flights, geotechnical device repair, and the radio setup. Additionally, the tools available were used in the Hab, Science Dome, GreenHab, and on EVA. Several days into the mission the RAM shopvac took up residence in the Hab due to the Hab Shopvac being inoperable.
In the Hab, a multitude of activities occurred. Downstairs the team prepared for EVAs and set up the Radio. The HSO is also a HAM radio operator (KO4HLC) she has made more than 50 contacts and completed the first known Morse Code contact from MDRS. She is currently working on updating the guidance for future radio operators. Upstairs was communal space. Here the team planned EVAs, filed reports and made delicious food.
Research updates
The team prepared multiple research projects to conduct over the course of the mission. See a summary of each in the sections below.
Geological-Geotechnical Surveys
The crew scientist proposed a series of geotechnical surveys to improve understanding of the geology near the Mars Desert Research Station and test the use of standard equipment in an analog environment. Over the course of 4 EVAs, 950 measurements were taken by the 2 geotechnical devices used to conduct the surveys. These measurements combined with samples taken for analysis and the aerial footage from the drone provides an extensive amount of information about the areas surveyed. There is a paper already under development related to the surveys during the mission.
Rock Sampling device test
The Pegasus Scoop was designed and built as a component for a larger robotic system. The crew scientist brought the scoop from Portugal for testing in a realistic environment. The scoop successfully performed its function. It will later be incorporated into a larger system.
Drone Payload testing
The drone brought by the executive officer was used for aerial footage and to conduct payload delivery tests while on EVA and to an EVA deploying from the Astronomy dome. Wind conditions limited the amount of testing completed. A total of 19 test flights were conducted with 17 being completed. The remaining 2 flights had to be aborted due to wind conditions mid-flight.
Passive watering tests
The passive watering system brought by the commander consisted of terracotta spikes with 500 mL plastic water bottles. On Sol 1 the 10 spikes were deployed into planters deep enough to utilize the spikes. The spikes were refilled when the bottles went dry, the spikes never went dry. This provided constant water flow directly to the plant’s roots. The remaining planters were assessed for their watering needs and watered as required. The total amount of watering done via both methods was collected.
Luxury crop sprouting test – Cacao
There were three cacao (chocolate) pods of different varieties of cacao brought to MDRS. They included 2 varieties of Trinitario (Red and Yellow) and Jaca a variant of Forastero. The seeds and seedlings require heat and humidity to sprout. On Sol 1 the three pods were opened, the seeds cleaned and set out for germination in trays on moist paper towels. On Sol 3 the seeds were rinsed and transferred to fresh paper towels. On Sol 5 the individual pots were prepared and the seeds showing signs of germination were planted. On Sol 10 multiple seeds showed signs they would be sprouting their cotyledons shortly. The first cotyledons were visible on Sol 12. Seeds were watered every day and the temperature and humidity monitored.
In Conclusion
The 12 Sol mission was a success from a research and outreach perspective. Team Pegasus looks forward to working together in the future and continuing the research started here.
Journalist Report – May 26th
Crew 281 Journalist Report 26May2023
Journalist Report
By: Rachel Jones
12th Sol, It’s Time for Crew 281 to Roll
My morning started very early. I wanted to see to the Sun rise on our last day of Mars. The Sun peaked the horizon at about 6:05 AM. As I reflected on this wonderful experience, I couldn’t be more grateful to have such amazing team mates.
Ana and Ritu were up early preparing for their longest EVA yet. Their explorative spirt and prospective on future scientific scanlines took them out to the Special Region (Burpee Quarry). First, Ritu conducted additional payload tests with the Medical Device Delivery Drone, “Peggy.”. Then, they traveled back to Marble Ritual to get additional aerial shots of the first scanline. Finally, they went out to the Special Region where dinosaur bones are being excavated. This EVA wrapped up all of Ana and Ritu’s scientific goals here at MDRS.
After MDRS, Ritu plans to publish her findings in a peer-reviewed journal. The data from her experience will be used for performing basic ratings and calculations for a larger drone system that will be capable of carrying a heavier load and of carrying out safe flights in high winds and lower air density.
Ana will further analyze her collected samples and the images that Ritu provided for further publication. Ultimately, Ana wants to publish a geotechnical zoning map for future MDRS crews.
KC was Hab Comm for the EVA. He then inventoried the entire kitchen. One of the things he wished he had was more preemptive information on kitchen supplies before he came to MDRS. He wants to give future crew better information. KC has been doing an amazing job as the Crew Engineer. Going above and beyond daily maintenance.
Megan made fudge brownies as a surprise for the EVA team while they were gone. Afterward, she spent almost six hours closing down the GreenHab for the season. For the finale, she brought in six kilograms of produce.
I, surprise, surprise, spent my day on the radio. I’ve learned a lot about HF operations. I made some mistakes. I learned how to jump into a pile-up (i.e. when multiple stations are trying to reach one station). It was really on this last day of operations that I got to experience some of the magic of amateur radio. Early in the morning, I got to hear operators from Australia and Japan. No, I didn’t get through those pile-ups, but it was amazing just to hear the other side of the world from a portable 20 Watt station in Utah.
As we close our journey here at MDRS, we think about the magic of what we are doing and why we are doing it. Each of us have a different dream we are pursuing. Ana, Ritu, and Megan want to be the first women on Mars. KC wants to farm on Mars. Ana wants to develop geotechnologies for building on Mars. I am passionate about space communications. Each of have our dreams to further human presence on Mars.
Thank you MDRS and thank you to all that have been following our journey. 73s.
Crew Photos – May 25th
Sol Summary – May 25th
Crew 281 Sol Summary Report 25 May 2023
Sol: 11
Summary Title: Geological miracles
Author’s name: Ritupriya Patil (XO) and Megan Kane (Commander)
Mission Status: Nominal
Sol Activity Summary:
EVA:
This was another long EVA with a crew of three. The crew entered the airlock exiting at 0753 hrs. They first checked all the rovers and collected the hours and SoC information, and it was verified that all were securely plugged in. They then unplugged Opportunity and Curiosity and headed out for the first destination at 0758 for the intersection of Cow Dung and Watney Road. The area seemed like a former riverbed; the crew saw lots of sedimentary rocks possibly washed by the flowing water in the past. They also witnessed calcite compressions on the sides of the pathway. The crew walked no more than 600 m and explored. Ana (Crew Scientist) collected 7 rocks (1167 g) and 1 Tube of soil (72 g). Ritu (XO, EVA Lead) deployed the drone to capture aerial.
For lunch the crew had leftovers and then worked on mission and research summaries. Rachel accomplished her goal of 50 contacts on her ham radio set up across 2 nations – USA and Canada. KC did his routine checks along with inventory work. Megan prepped green hab for end of mission dismantling. Ana did the same for science dome and the RAM. Ritu worked on the daily reports and EVA planning for tomorrow, which will be the longest of all.
For dinner Megan prepared potato vegetable soup for dinner and then crew jumped on a call with the current Hi-SEAs crew called IKE-LOA. During the call, both crews exchanged their experience with research, facilities and showed around their respective habitats. It was a great way to connect with another crew doing similar mission, Hi-SEAs was on the Moon and MDRS crew was on Mars. This was also a great opportunity to network with other space explorers.
Look Ahead Plan: Tomorrow we will do 2-person, long range EVA to run some more drone payload tests. Next we will stop at Marble Ritual for images of scanlines that were measured on earlier EVAs. Last stop will be the Special Region in the north.
Anomalies in work: None
Weather: Winds
Crew Physical Status: Nominal
EVA: Completed 1 today
Reports to be filed:
Greenhab Report
EVA Report
EVA Request
Ops Report
Journalist Report
PHOTOS!
Photo of the Day
Support Requested:
Mission support is requested in case of storms if the antenna needs to come down or be retrieved.
Operations Report – May 25th
Crew 281 Operations Report 25 May 2023
SOL: 11
Name of person filing report: Megan Kane (Commander), Ritupriya Patil (Executive Officer) and KC Shasteen (Crew Engineer)
Non-nominal systems:
Electrical outlets in kitchen
Remote observatory
EVA Helmet
Notes on non-nominal systems: None
ROVERS
Spirit rover used: No
Hours: (before EVA): 222.5
Hours: (after EVA): NA
Beginning charge: (Before EVA): 100%
Ending charge: (On return from EVA, before recharging): NA
Currently charging: Yes
Opportunity rover used: Yes
Hours: (before EVA): 131.2
Hours: (after EVA): 132.1
Beginning charge: (Before EVA): 100%
Ending charge: (On return from EVA, before recharging): 95%
Currently charging: Yes
Curiosity rover used: Yes
Hours: (before EVA): 240.2
Hours: (after EVA): 241.0
Beginning charge: (Before EVA): 100%
Ending charge: (On return from EVA, before recharging): 75%
Currently charging: Yes
Perseverance rover used: No
Hours: (before EVA): 264.6
Hours: (after EVA): NA
Beginning charge: (Before EVA): 100%
Ending charge: (On return from EVA, before recharging): NA
Currently charging: Yes
General notes on rovers: Oppy and Curie used on long EVA to the intersection of Cow Dung and Hwy 1101 and back. Operation was nominal.
Summary of Hab operations:
WATER USE: 32.4 gallons
Water (static tank): 202.2 gallons
Static tank pipe heater (on or off): off
Static tank heater (On or off): off
Toilet tank emptied: no
Summary of internet:
nominal
Summary of suits and radios: All suits and radios are fully recharged.
Summary of GreenHab operations:
WATER USE: 2.25 gallons
Heater: Off
Supplemental light: Off
Harvest: in grams : None
Summary of Science Dome operations:
Dual split: Heat and AC, Off
Cacao Experiment:
Watered cacao seedlings
Initial cleaning in preparation for exiting the station.
Summary of RAM operations: (Tools used, work done)
Initial cleaning completed in preparation for exiting the station.
Summary of any observatory issues: NA
Summary of health and safety issues:
No new issues to report.
Questions, concerns and requests to Mission Support:
Request Mission Support assistance to take down or retrieve the ham radio antenna in case of emergency or high winds.
You must be logged in to post a comment.